Table 7-3
Alton Company had the following activity in its inventory account during April 20X9.

Cost per
Date Activity Units Unit Cost Total
April 1 Beginning inventory 100 $3.00 $300
April 3 Purchase 40 3.10 124
April 7 Sale 50
April 12 Purchase 50 3.20 160
April 16 Sale 70
April 23 Sale 40
April 30 Purchase 60 3.30 198

Units in beginning inventory 100 units
Units purchased 150 units
Units sold 160 units


Referring to Table 7-3, what is the ending inventory balance at April 30, 20X9, for Alton Company if the company uses perpetual FIFO as its inventory valuation method?

$297.50

$358.00

$270.00

$198.00

$294.00
Answer #1

Under Perpetual method, the records are updated after each transaction.

FIFO method states that inventory purchased first should be sold first

Units in ending inventory = 100+40-50+50-70-40+60 = 90 units

Hence, ending Inventory balance = 30*3.20 + 60*3.30

= $294

Hence, the answer is $294.00